OpenStudy (moonlight93):

The rate constant of a reaction is 7.0 x10^-3 s^-1 at 25 degree C, and the activation energy is 33.6 kj/mol. What is k at 75 degree C? I know I have to use ln(k1/k2) = (Ea/R)[(1/T2)-(1/T1)] .. but I'm still confuse. Any help?

OpenStudy (anonymous):

ln (k1/k2) = Ea/R (1/T2-1/T1) R is the gas constant with Joule units: 8.314 J/mol-K You will need to switch your Activation Energy into Joule units: 33600 J and your Temperatures into Kelvin. Once you have done this, plug in and solve for k2. Variables: k1 = 4.7 x 10^(-3) k2 = ? T1 = 289 K T2 = 348 K Ea = 33600 J R = 8.314 J/mol-K